Grade 12Physical Chemistry

Assign oxidation number to the underlined elements in each of the following species: (a) NaH2PO4 (b) NaHSO4 (c) H4P2O7 (d) K2MnO4 (e) CaO2 (f) NaBH4 (g) H2S2O7 (h) KAl(SO4)2.12 H2O ?

sudhanshu
12 Years agoGrade 12

1 Answer

Gaurav
11 Years ago
Hello Student
(a) NaH2PO4 :- Oxidation number of P in NaH2PO4 is 1(+1)+2(+1)+1(x)+4(-2)=0 = X= +5
(b) NaHSO4 :- Oxidation number of S in NaHSO4 is 1(+1)+1(+1)+x+4(-2)=0 = x = +6
(c) H4P2O7: - Oxidation number of P in H4P2O7 is 4(+1)+2(x)+7(-2)=0 = x = +5
(d) K2MnO4:- Oxidation number of Mn in K2MnO4 is 2(+1)+1(x)+4(-2)=0 = X = +7
(e) CaO2 :- Oxidation number of O in CaO2 is +2+2x =0 = x = -1
(f) NaBH4:- Oxidation number of B in NaBH4 is 1(+1) + x + 4(-1) = 0 = x = +3
(g) H2S2O7 :- Oxidation number of S in H2S2O7 is 2(+1)+2(x)+7(-2)= 0 = x = +6
(h) KAl(SO4)2.12 H2O:- Oxidation number of S in KAl(SO4)2.12 H2O is +1+3+2x-16=0 = x = +6